OpenStudy (anonymous):

d/dx x^2-x-1=2x-1? or = 2x-x-1 does it have to have a parenthesis?

OpenStudy (amistre64):

-1 is a constant to begin with; or rather, -1x^0

OpenStudy (amistre64):

to derive it, follow the power rule

OpenStudy (anonymous):

d/dx (x^2-x-1) or d/dx (x^2-x) -1

OpenStudy (amistre64):

what does the "have to have ..." part mean?

OpenStudy (anonymous):

the problem says \[d/dx x^2-x-1 \] is it the same as \[d/dx(x^2-x-1)\]??

OpenStudy (anonymous):

first one is not clear

OpenStudy (amistre64):

\[\frac{d}{dx}(1x^2 -x -1)\] \[\frac{d}{dx}(1x^2)- \frac{d}{dx}(1x^1)- \frac{d}{dx}(1x^0)\] power rule says to bring down the exponent and subtract 1 \[\frac{dx}{dx}2*1x^{2-1}- \frac{dx}{dx}(1*1x^{1-1})- \frac{dx}{dx}(0*1x^{0-1})\] \[\frac{dx}{dx}2x^{1}- \frac{dx}{dx}(1x^{0})- \frac{dx}{dx}(0x^{-1})\] and dx/dx=1 \[2x-1\]